# Neural Scaling Laws on Graphs

### v1

@@ 5.2 The Effect of Model Depth
Another unique question in the graph domain is whether the model depth would affect the model scaling law, or more specifically, the scaling law coefficients in Equation 3. Kaplan et al. (2020) find that in the NLP domain, the model depth does not have effects on the scaling coefficients, i.e., the scaling curves for transformers of different numbers of model layers will overlap. Therefore, conducting preliminary experiments to determine the optimal number of layers for large language models (LLMs) is not required, which effectively narrows the hyperparameter search space. Since the deep graph models have different architectures from vanilla transformers, it is desired to study whether the above conclusion still holds.
(a) ogbg-molhiv
(b) reddit-multi-12k
Figure 8: Performance curves of GIN with varying sizes trained on two subsets. The simple subset and the complex subset have the same number of graphs but different numbers of edges. The shade corresponds to the performance variance of repeated experiments.
Figure 9: Performance curves of GIN of varying sizes trained on two subsets. The simple subset and the complex subset have the same number of edges but different numbers of graphs. The performance differences between the two curves are much smaller than the curves in Figure 8(a), suggesting that the number of edges is a better data metric than the number of graphs.
To ensure the generalization of our investigations, we choose two representative GNNs (GCN & GIN) and two powerful graph transformers (SAT & GraphGPS). For these models, we choose the layer number in $[3,4,5]$ and fix it but change the model width to attain varying model sizes. All the models are trained on the whole ogbg-ppa training set and then evaluated for model performance. The model scaling curves of models with different model depths are presented in Figure 6 and 7. Observations can be made as follows: (1) The scaling behaviors of models with different numbers of layers are distinct. For each model, the shape of the scaling law curves would change when the model depth changes, exhibiting different increasing speeds and upper bounds for performance. (2) The optimal number of layers for different models could be different. For GCN and GraphGPS, models with $3$ layers have the best performance, but for GIN the optimal layer number is $4$, and for SAT is $5$. Similar observations on more datasets can be found in Appendix D.
@@
Observation 3. Deep graph models with varying depths will have different model scaling behaviors.
Current observations support that there is no one-size-fits-all model depth for existing graph learning models. The optimal number of model layers could be task-specific and model-specific. We emphasize the necessity of preliminary experiments for the optimal number of model layers before scaling up the model parameters. Notably, we do not have any principle guidance helping to automatically select the best model layer so far. Such mystery could be attributed to the complicated non-parametric aggregation design in different GNNs with various expressiveness (Anonymous, 2024) and inductive bias (Battaglia et al., 2018). Therefore, it is likely that there is no principle across models. Instead, there could be some rules specific to tasks and models without a consensus. Therefore, such analysis is out of scope targeting a general principle across different model architectures and different tasks. We leave the deeper investigation on it as the future work.
6 Reforming Data Scaling Law on Graphs
@@
In this section, we address the irregular-size problems of graphs for the data scaling law. Specifically, we re-examine data metrics in previous studies and reveal the disadvantage of the number of graphs as the data metric in the data scaling law. Instead, we identify the total number of edges in the training set as a better metric with experimental evidence. Furthermore, the data scaling law with the number of edges as the data metric can be naturally extended to node classification and link prediction tasks, potentially providing a unified view of data scaling behaviors in the graph domain.
(a) Node classification on ogbn-arxiv, $R^{2}=0.97$.
(b) Link prediction on ogbl-collab, $R^{2}=0.99$.
@@
6.1 Number of Edges as Data Metric
In previous sections, the data metric in scaling laws is the subsampling ratio of total graphs, i.e., how many graphs in the whole training set are used (Rosenfeld et al., 2019; Chen et al., 2023). Nonetheless, using the number of graphs as the metric neglects the size irregularity of graphs. For instance, a graph with $10$ edges and a graph with $1,000$ edges are treated equivalently when the number of graphs is the metric, even though they could differ remarkably in training costs and impacts on model performance improvement.
To illustrate the disadvantages of adopting the number of graphs as the data metric, we first divide the training sets of one dataset into two subsets as follows: (1) We first arrange all graph samples of class $i$ into a list $\mathcal{G}_{i}$ in an ascending order of their numbers of edges. (2) We choose the first half of all $\mathcal{G}_{i}$ as the “simple” training subset, and then the second half as the “complex” subset. Thus, these two subsets have the same number of graphs while the graphs in the complex subset have more edges. The validation and test sets remain as the original. Then we train GIN of varying sizes on these two subsets and report the performance on ogbg-molhiv and reddit-multi-12k in Figure 8. Results on more datasets can be found in Appendix E.
If the number of graphs can accurately metric the data volume, the curves of simple and complex subsets in Figure 8 should have small differences or even overlap with each other. However, their performance curves exhibit considerable differences, indicating that the two subsets contain different amounts of data. For comparison, we attain the two subsets in another way: the simple subset has the first $m$ samples in $\mathcal{G}_{i}$, and the complex subset has the remaining $n$ samples, where the two subsets contain the same number of edges. It is easy to validate that $m>n$. Then we train models of varying sizes on these two subsets and report their performance on ogbg-molhiv in Figure 9. Compared to Figure 8(a), the curves in Figure 9 have much smaller differences, which indicates that the total number of edges could be a better metric than the number of graphs.
Observation 4. The number of edges is a better data metric for data scaling law compared to the number of graphs.
In addition, using the total number of edges as the data metric in scaling laws has two more advantages: (1) Knowing the GNN architecture and the number of edges in a dataset, the compute cost of training can be directly calculated, which enables us to build connections between data scaling law and compute scaling law as in Kaplan et al. (2020). The detailed explanation is in Appendix F. (2) Furthermore, using the number of edges as the data metric potentially provides a unified way to describe the data scaling behaviors of node classification and link prediction tasks, which will be thoroughly studied in the next subsection.
6.2 Results on Node Classification and Link Prediction
There is only one graph in most of the node classification and link prediction datasets, thus it is infeasible to use the number of graphs to measure the training data volume. Nonetheless, the total number of edges is able to serve as the data metric for these tasks. Moreover, with the total number of edges as the data metric, we can potentially attain a unified data scaling law on graphs by directly extending Eq. 3 to node classification and link prediction tasks. To verify this, we conduct experiments on the data scaling behaviors on both node classification and link prediction tasks. For node classification, we train a $3$-layer GraphSAGE (Hamilton et al., 2017) on ogbn-arxiv, where the hidden vector size is set to $128$. For link prediction, we train a $3$-layer SEAL (Zhang et al., 2021) on ogbl-collab, where the model takes one-hop neighbors of the target node pairs as input and its hidden vector size is set to $256$. We take the total number of edges on which models are trained as the training set size for both tasks. The results are demonstrated in Figure 10. Similar to graph classification tasks, Eq. 3 fits well to the data scaling behaviors, as evidenced by the high $R^{2}$ values.
Observation 5. Node classification, link prediction, and graph classification tasks could follow the same form of data scaling law with the number of edges as data metric.
@@ 7 Conclusions
In this work, we take the first step to investigate neural scaling laws on graphs. In particular, we focus on two basic forms of neural scaling laws: the model scaling law and the data scaling law. We first verify the formulations that can describe both scaling behaviors in the graph domain. For the model scaling law, we identify and investigate some graph-specific phenomena, revealing different observations from CV and NLP domains. For the data scaling law, we reform it by adopting the number of edges as the data metric and extend the reformed version to node classification and link prediction tasks. Our findings shed light on the scaling behaviors of deep graph models and can provide valuable instructions for developing large graph models.
Limitations. Next, we discuss the limitations of this work, which could be further explored as future works. (1) There is no large graph dataset of similar sizes as the large CV or NLP datasets. We conduct our experiments on the largest academic graph datasets that we can attain. (2) Since there has not been a well-established pretraining task for graph learning, we limit our investigation to the supervised learning on graphs. (3) In our exploration, we assume that all node (edge) features are in a unified feature space. We do not consider heterogeneous feature problems in this work.
Impact Statements
@@
Datasets. The basic situations of the datasets we use are listed in the Tabel 1 and  2 below.
@@
Table 2: Node Classification and Link Prediction Datasets Statistics.
ogbn-arxiv ogbl-collab
Domains Citation Social
#Nodes 169,343 235,868
#Edges 1,166,243 1,285,465
Tasks Node Classification Link Prediction
Metrics Accuracy Hits@20
Training Details. For GIN and GCN models, we set the initial learning rate to 0.001 and use Adam optimizer (Kingma & Ba, 2014) in the training process. For SAT and GraphGPS, we directly adopt the hyperparameters and optimizers that the authors present in their GitHub repos. For ogbg-ppa, ogbn-arxiv, and ogbl-collab datasets, the training epoch number is 50. For pcqm4mv2, ogbg-molhiv, ogbg-molpcba, and reddit-multi-12k datasets, the training epoch number is 100. We choose the test performance with the best corresponding validation performance as the final performance of the model.
@@ Appendix B More Results on the Validity of the Neural Scaling Law on Graphs
Figure 12: The model and data scaling behaviors of GCN on ogbg-ppa. For model scaling, the color corresponds to the training data fraction; for data scaling, the color corresponds to the number of model parameters. The higher performance is, the better model performance is. The high $R^{2}$ values indicate that the scaling law curves fit well with the empirical values.
@@
Appendix C More Results on the Model Scaling Collapse
@@
Appendix D More Results on the Impacts of Model Depths
@@
Appendix E More Results of the Disadvantage of the number of Graphs as Data Scaling Metric
Here we present more results of the experiments in Section 6.1. The settings are the same as in Section 6.1, except we replace GIN with GCN. From the below figure, we observe that performance curves have obvious differences when two subsets have the same number of graphs. But when the two subsets have the same number of edges, the differences between the two curves become much smaller, supporting that the number of edges can better measure the data volume.
Figure 20: Performance curves of GCN of varying sizes trained on two subsets. In (a), the simple subset and the complex subset have the same graph number but different edge numbers. In (b), the simple subset and the complex subset have the same edge number but different graph numbers. The shade corresponds to the performance variance of repeated experiments. This figure supports the same conclusion as in Section 6.1.
Appendix F The Relationship between Edge Number and FLOPs
Here we show that the training compute can be calculated when we know the total edge number and the model architecture. We consider the general Message Passing Neural Networks, which can be formulated as above:

### v3

@@ 5 The Effect of Model Depth on Graph Scaling Laws
A unique question in the graph domain is whether the model depth would affect the model scaling law, or more specifically, the scaling law coefficients in Equation 3. Previous research [2] in the NLP domain finds that the model depth does not have effects on the scaling coefficients, i.e., the scaling curves for transformers of different numbers of model layers will overlap. Therefore, conducting preliminary experiments to determine the optimal number of layers for large language models (LLMs) is not required, which effectively narrows the hyperparameter search space.
@@
Figure 4: Model scaling behaviors of GNNs and graph transformers with varying model depths. The color corresponds to the model layer. All the models exhibit distinct scaling behaviors with different model depths.
Nonetheless, the deep graph models have different architectures, i.e., non-parametric aggregation layer, from vanilla transformers, which induces unique problems in graph learning such as over-smoothing [31] and over-squashing [32]. Thus, it is desired to study whether the conclusion from the NLP domain still holds on graphs.
Figure 5: Model scaling behaviors when fixing the number of aggregation layers or feed-forward layers.
To ensure the generalization of our investigations, we choose two representative GNNs (GCN & GIN) and two powerful graph transformers (SAT & GraphGPS). For these models, we choose the layer number in $[3,4,5]$ and fix it but change the model width to attain varying model sizes. All the models are trained on the whole ogbg-ppa training set and then evaluated for model performance. The model scaling curves of models with different model depths are presented in Figure 4. Observations can be made as follows: (1) The scaling behaviors of models with different numbers of layers are distinct. For each model, the shape of the scaling law curves would change when the model depth changes, exhibiting different increasing speeds and upper bounds for performance. (2) The optimal number of layers for different models could be different. For GCN and GraphGPS, models with $3$ layers have the best performance, but for GIN the optimal layer number is $4$, and for SAT is $5$. Similar observations on more datasets can be found in Appendix E.
Observation 2. Deep graph models with varying depths will have different model scaling behaviors.
Current observations support that there is no one-size-fits-all model depth for existing graph learning models. The optimal number of model layers could be task-specific and model-specific. We emphasize the necessity of preliminary experiments for the optimal number of model layers before scaling up the model parameters. Further, we look deep into the relationship between the non-parametric aggregation operation and the layer difference of graph model scaling. To achieve this goal, we first decouple the GIN layer into one feed-forward layer and one non-parametric aggregation layer, then repeat the above experiments by varying the number of feed-forward or aggregation layers while fixing the other to be 5. The results are shown in Figure 5. We observe that no matter whether fixing the aggregation layer or feed-forward layer, there will always be layer differences in model scaling. These observations suggest that the non-parametric aggregation could be one of the reasons for the layer difference. As for the observations when increasing the feed-forward layers, we speculate the local reception field of GNN (different from the global attention of vallina transformer) to be the reason. Notably, we do not have any principle guidance helping to automatically select the best model layer so far and we still need to select the optimal layer number for specific graph learning tasks. Thus, building a new model whose model scaling behaviors are agnostic to the number of layers could be beneficial to the development of large graph models.
6 Proper Data Metric of Graph Scaling Laws
@@
In this section, we look into the data scaling law on graphs. Specifically, we re-examine data metrics in previous studies and reveal the disadvantage of the number of graphs as the data metric in the data scaling law. For instance, a small molecule graph could have a large size difference compared with a social network. Instead, we identify the total number of nodes or edges in the training set as a better metric with experimental evidence.
In previous studies, the data metric in scaling laws is the number of graph samples, i.e., how many graphs in the whole training set [1, 22]. Nonetheless, using the number of graphs as the metric neglects the size variability of graphs. For instance, a graph with $10$ edges and a graph with $1,000$ edges are treated equivalently when the number of graphs is the metric, even though they could differ remarkably in training costs and impacts on model performance improvement. Empirically, we conduct two controlled experiments to show that the number of nodes or edges is the better data metric in general.
Figure 6: Comparison between the number of graphs and edges as data metric
Experiment 1: From the data scaling perspective
In this experiment, we show that the number of nodes (edges) as the data metric can provide a better fitting of scaling laws than the number of graphs. To highlight the contrast, the training set is scaled up in the following way: each time we add a subset with the same number of graphs but an increasing number of total nodes/edges. However, the graph size distribution often varies among classes of real-world datasets. To eliminate the bias, we adopt the setup in  [33] and construct a synthetic dataset with a motif recognition task. Notably, in the setup [33], the synthetic graphs are generated as Barabasi Albert graphs, whose nodes are connected with a fixed number of edges. Thus, we can view the numbers of nodes and edges as equivalent in this setting. Specifically, the dataset includes three classes, i.e., three motifs (Cycle, Grid, House) to recognize. We generate 8 subgroups for each class, where each subgroup contains 100 graphs with the same number of nodes $i$ ($i\in[25,50,100,200,400,800,1600,3200]$). Each time to increase the training set size, we add the subgroups of three classes with the same number of nodes $i$ ($i$ increases monotonically from 25 to 3200). Then we can ensure that the total number of graphs from each class is balanced. After recording the model performance of varying training set sizes, we fit the data scaling law to the empirical values with the number of nodes (edges) as the data metric repectively. The results of the number of edges are shown in Figure 6. Results with the number of nodes as the metric can be found in Appendix G. From the figure, we observe that the number of nodes as the data metric gives a more accurate fitting with a higher $R^{2}$ value, which supports our argument.
Experiment 2: From the model scaling perspective
For real-world datasets, we contrast the two metrics by comparing the model scaling curves of the same model on two training sets with the same number of graphs but different number of nodes/edges. Since this study focused on in-domain graph learning, the graph samples tend to have similar average degrees, as shown in Figure 7. The degree distributions of other graph datasets have the same trend, which can be found in Appendix F. Hence, we could treat the number of nodes and edges as nearly equivalent metrics. We discuss the situations with the number of edges as the metric in this subsection and results with the number of nodes can be found in Appendix G. We first divide the whole training sets of one dataset into two subsets as follows: (1) We first arrange all graph samples of class $i$ into a list $\mathcal{G}_{i}$ in an ascending order of their numbers of edges. (2) We choose the first half of all $\mathcal{G}_{i}$ as the “simple” training subset, and then the second half as the “complex” subset. Thus, these two subsets have the same number of graphs while the graphs in the complex subset have more edges. The validation and test sets remain as the original. Then we train GIN of varying sizes on these two subsets and report the performance on ogbg-molhiv in Figure 8.
Figure 7: Average degree distribution of ogbg-molhiv dataset.
Figure 8: Model scaling curves of GIN on two training subsets of different sizes. (a) Two subsets have the same total number of graphs but different total number of edges. (b) Two subsets have the same total number of edges but different total number of graphs. The shade corresponds to the performance variance of repeated experiments.
If the number of graphs can accurately metric the data volume, the curves of simple and complex subsets in Figure 8(a) should have small differences or even overlap with each other. However, their performance curves exhibit considerable differences, indicating that the two subsets contain different amounts of data. For comparison, we attain the two subsets in another way: the simple subset has the first $m$ samples in $\mathcal{G}_{i}$, and the complex subset has the remaining $n$ samples, where the two subsets contain the same number of edges. It is easy to validate that $m>n$. Then we train models of varying sizes on these two subsets and report their performance on ogbg-molhiv in Figure 8(a).
Compared to Figure 8(a), the curves in Figure 8(b) have much smaller differences, which indicates that the total number of edges could be a better metric than the number of graphs.
Observation 3. The number of edges is a better data metric for data scaling law compared to the number of graphs.
In addition, using the total number of edges as the data metric in scaling laws has two more advantages: (1) Knowing the GNN architecture and the number of edges in a dataset, the compute cost of training can be directly calculated, which enables us to build connections between data scaling law and compute scaling law as in [2]. The detailed explanation is in Appendix I. (2) Furthermore, using the number of edges as the data metric potentially provides a unified way to describe the data scaling behaviors of node classification and link prediction tasks, which will be thoroughly studied in the following section.
7 Results on Node Classification and Link Prediction
Furthermore, the data scaling law with the number of nodes (edges) as the data metric can be naturally extended to node classification and link prediction tasks, potentially providing a unified view of data scaling behaviors in the graph domain. There is only one graph in most of the node classification and link prediction datasets, thus it is infeasible to use the number of graphs to measure the training data volume. Nonetheless, the total number of nodes (edges) is able to serve as the data metric for these tasks. Moreover, with the total number of nodes (edges) as the data metric, we can potentially attain a unified data scaling law on graphs by directly extending Equation 3 to node classification and link prediction tasks. To verify this, we conduct experiments on the data scaling behaviors on both node classification and link prediction tasks. Here we only provide results with the number of edges as metric, for the results with the number of edges as metric, please refer to Appendix H. For node classification, we train a $3$-layer GraphSAGE [34] on ogbn-arxiv, ogbn-products and ogbn-proteins, where the hidden vector size is set to $128$. For link prediction, we train a $3$-layer SEAL [35] on ogbl-collab, ogbl-ppa and ogbl-citation2, where the model takes one-hop neighbors of the target node pairs as input and its hidden vector size is set to $256$. We take the total number of edges on which models are trained as the training set size for both tasks. The results are demonstrated in Figure 9. Similar to graph classification tasks, Equation 3 fits well to the data scaling behaviors, as evidenced by the high $R^{2}$ values.
Observation 5. Node classification, link prediction, and graph classification tasks could follow the same form of data scaling law with the number of nodes (edges) as data metric.
@@ 8 Conclusions
In this work, we take the first step to investigate neural scaling laws on graphs. In particular, we focus on two basic forms of neural scaling laws: the model scaling law and the data scaling law. We first verify the formulations that can describe both scaling behaviors in the graph domain. For the model scaling law, we identify and investigate some graph-specific phenomena, revealing different observations from CV and NLP domains. For the data scaling law, we reform it by adopting the number of edges as the data metric and extend the reformed version to node classification and link prediction tasks. With more attention has been attracted to the graph foundation models [36, 37, 38, 39], our findings shed light on the scaling behaviors of deep graph models and can provide valuable instructions for developing large graph models.
Limitations. Next, we discuss the limitations of this work, which could be further explored as future works. (1) There is no large graph dataset as large as CV or NLP datasets. We conduct our experiments on the largest graph datasets that we can attain in academics. Even though, this initial study is still meaningful. Only after the data scaling law on graphs has been verified, more efforts would be made to construct larger graph datasets, and the experience from this work could also be used to guide the process. For instance, Tang et al [40] adapt the data scaling scaling perspective and improve the crosse-domain graph learning performance by generating synthetic graph data for training by diffusion. (2) Since there has not been a well-established pretraining task for graph learning, we limit our investigation to the supervised learning on graphs. (3) In our exploration, we assume that all node (edge) features are in a unified feature space. We do not consider heterogeneous feature problems in this work.
@@
Datasets. The basic situations of the datasets we use are listed in the Table 12 and  3 below.
@@
Table 2: Node Classification Datasets Statistics.
ogbn-arxiv ogbn-products ogbn-proteins
Domains Citation E-commerce Biology
#Nodes 169,343 2,449,029 132,534
#Edges 1,166,243 61,859,140 39,561,252
Metrics Accuracy Accuracy ROC-AUC
Table 3: Link Classification Datasets Statistics.
ogbl-ppa ogbl-collab ogbl-citation2
Domains Citation E-commerce Biology
#Nodes 576,289 235,868 2,927,963
#Edges 30,326,273 1,285,465 30,561,187
Metrics Hits@100 Hist@20 MRR
Training Details. For GIN and GCN models, we set the initial learning rate to 0.001 and use Adam optimizer [41] in the training process. For SAT and GraphGPS, we directly adopt the hyperparameters and optimizers that the authors present in their GitHub repos. For all the models, we opt against the learning rate with a scheduler. For ogbg-ppa, ogbn-arxiv, ogbn-products, ogbn-proteins, ogbl-ppa, ogbl-citation2 and ogbl-collab datasets, the training epoch number is 50. For pcqm4mv2, ogbg-molhiv, ogbg-molpcba, and reddit-multi-12k datasets, the training epoch number is 100. We choose the test performance with the best corresponding validation performance as the final performance of the model.
@@ Appendix B More Results on the Validity of the Neural Scaling Law on Graphs
Figure 11: The model and data scaling behaviors of GCN on ogbg-ppa. For model scaling, the color corresponds to the training data fraction; for data scaling, the color corresponds to the number of model parameters. The higher the performance is, the better the model performance is. The high $R^{2}$ values indicate that the scaling law curves fit well with the empirical values.
@@
Appendix C Model Scaling Law on Limited Graph Data
(a) Model scaling on ogbg-molpcba
(b) Training losses of models
(c) Validation losses of models
Figure 13: (a) The empirical performance of larger models falls below the prediction of model scaling law. (b)(c) The training losses and validation losses of models with varying sizes. The color corresponds to the number of model parameters. The model with larger sizes suffers more serious overfitting.
The model scaling collapse poses difficulty in building large graph models following the model scaling law. Thus, it is essential to explore why the collapse happens and how to mitigate it. In this work, we conjecture that one of the potential reasons for the scaling law collapse is overfitting. To dive deeper into the problem, Figure 13(b) and 13(c) present the training loss and validation loss curves, respectively. All the training losses in Figure 13(b) keep dropping in the training process. The larger the model is, the smaller the corresponding training loss will be. Nonetheless, the validation losses of models with more than $10^{7}$ parameters start increasing after several epochs as shown in Figure 13(c), which indicates overfitting. Moreover, the models whose performance falls behind the scaling law predictions suffer more serious overfitting. For instance, models of 7.4M, 13M, and 20M parameters overfit the most in Figure 13(c), and they are the three models on which model scaling law fails in Figure 13(a). Based on these observations, overfitting can be one potential reason for the model scaling collapse. For more results between the model size and overfitting, please refer to Appendix D.
To eliminate the effect of model scaling collapse, it is crucial to overcome the overfitting. [2] suggests that overfitting would happen if the model size is too large for the training set. To investigate whether more training data can mitigate overfitting, we train a model of fixed number parameters on different fractions of the training set and illustrate the validation loss curves in Figure 14. The model trained on $10$ percent of the original ogbg-molpcba dataset suffers serious overfitting as the validation loss curve suggests. However, the overfitting is significantly mitigated when the data fraction is raised to $50$ percent, which indicates the effectiveness of enlarging the training set. Another evidence is that we have not observed overfitting or model scaling collapse in models of similar sizes ($10^{6}$ parameters) trained on larger datasets, i.e., pcqm4mv2 and ogbg-ppa in Section 4.
Appendix D More Results on the Model Scaling Collapse
@@
Appendix E More Results on the Impacts of Model Depths
@@
Appendix F The Average Degree Distributions of More Datasets
Here we present the average degree distributions of more datasets in Figure 21. The samples in one dataset tend to have similar average degrees, further supporting our argument in Section 6.
(a) bace
(b) bbbp
(c) molchembl
(d) molpcba
(e) Ogbg-ppa
(f) tox21
Figure 21: The average degree distribution of more datasets.
Appendix G Results with the Number of Nodes as metric for Section 6
Here we present more results of the comparison of the data metric of the data scaling law. Figure 22 shows that the number of nodes as metric could also give better fitting than the number of graphs, supporting the conclusion of Experiment 1 in Section 6. Figure 23 shows that training subsets with the same number of total nodes but different numbers of graphs tend to have similar scaling behaviors, supporting the conclusion of Experiment 2 in Section 6.
Figure 22: Comparison between the number of graphs and the number of nodes as the data metric
Figure 23: Model scaling curves of GIN on two training subsets of different sizes. The shade corresponds to the performance variance of repeated experiments.
Appendix H The Data Scaling Law of Node Classification and Link Prediction with the Number of Nodes as Metric
Here we present the results of the data scaling law on node classification and link prediction tasks with the number of nodes as the data metric. From Figure 24, we find that the number of nodes also gives good fitting results as the data metric.
Figure 24: Both node classification and link prediction follow the data scaling law with the total node number as the data metric.
Appendix I How do Our Findings Connect to Compute Scaling and Chinchilla Scaling Law
Here we first show that the training compute can be calculated when we know the total edge number and the model architecture. We consider the general Message Passing Neural Networks, which can be formulated as above: